OpenStudy (anonymous):

17. ∠a and ∠b are supplementary. If m∠a = 2x + 1 and m∠b = 3x - 6, what is the measure of ∠b? A.35° B.37° C.71° D.105°

OpenStudy (anonymous):

supplementary angles are two angles, when added, equal 180 degrees. so if <a and <b are supplementary, then they equal 180 degrees. 2x + 1 + 3x - 6 = 180 -- combine like terms 5x - 5 = 180 -- add 5 to both sides 5x = 185 -- divide both sides by 5 x = 37 <b = 3x - 6 <b = 3(37) - 6 <b = 111 - 6 <b = 105 degrees any questions ?